OpenStudy (anonymous):

If the sin 30 is 1/2, then the cos and it has a blank an equal sign and another blank

zepdrix (zepdrix):

Sine and Cosine are `co-functions`. \[\Large\rm \sin(\theta)=\cos(90-\theta)\] So if \(\Large\rm \sin(30)=\frac{1}{2}\) Then, \(\Large\rm \cos(90-30)=?\)

OpenStudy (anonymous):

so 60

OpenStudy (anonymous):

???

OpenStudy (anonymous):

not to butt in , but this is the second time i have seen a question that starts with " If the sin 30 is 1/2" like saying "if two plus two is four, then ..." it makes no sense

zepdrix (zepdrix):

Yah they didn't really give enough information to answer the question :) But I could see where it was headed lol

zepdrix (zepdrix):

That is pretty frustrating.

OpenStudy (anonymous):

you can fill it in with literally anything \[\cos(30)=\frac{\sqrt2}{2}\] or \[\cos(60)=\frac{1}{2}\] or whatever you like

OpenStudy (anonymous):

lol can someone just help me then please

zepdrix (zepdrix):

@Samidancer16 Yes,\[\Large\rm \cos(90-30)=\cos(60)\]But relate that back to this:\[\Large\rm \color{royalblue}{\sin(\theta)=\cos(90-\theta)}\] If we plug in 30 for our angle we get this relationship:\[\Large\rm \color{royalblue}{\sin(30)=\cos(60)}\]

zepdrix (zepdrix):

What does that tell you about cos(60)?

OpenStudy (anonymous):

so the measurement is 60 for an angle?

OpenStudy (anonymous):

Im sorry im bad at this

zepdrix (zepdrix):

Just pay attention to the equals sign. He's telling us a lot.

zepdrix (zepdrix):

sin(30) is 1/2, they told us that. What does that tell us about what cos(60) is equal to? :p

OpenStudy (anonymous):

it is not your fault the question is ambiguous and clearly written by a moron

zepdrix (zepdrix):

\[\Large\rm \color{#CC0033}{\sin(30)}=\cos(60)\]Mmk :( ugh We'll replace our sin(30) with 1/2, since they are equivalent.\[\Large\rm \color{#CC0033}{\frac{1}{2}}=\cos(60)\]Understand what I did? sin(30) is 1/2. So we rewrote sin(30) as 1/2.

zepdrix (zepdrix):

That's what the relationship is telling us.

OpenStudy (anonymous):

i dont understand where the60 and the 1/2 is coming from

zepdrix (zepdrix):

Sine and Cosine are co-functions. This is just a formula that you'll have to accept for now :)\[\Large\rm \sin(\theta)=\cos(90-\theta)\]The angle we care about is \(\Large\rm \theta=30\) Plugging that value in is what gives us our relationship.\[\Large\rm \sin(30)=\cos(90-30)\]90-30=60, yes?\[\Large\rm \sin(30)=\cos(60)\]

zepdrix (zepdrix):

They told us that:\[\Large\rm \sin(30)=\frac{1}{2}\]That was given to us.

zepdrix (zepdrix):

So if:\[\Large\rm \sin(30)=\frac{1}{2}\]and\[\Large\rm \sin(30)=\cos(60)\]What can we say about cos(60)?

OpenStudy (anonymous):

okay i get that now where is the 1/2 coming from

zepdrix (zepdrix):

ya +_+ weird stuff, ik ik

OpenStudy (anonymous):

okay okay nvm

OpenStudy (anonymous):

so is the answer 60; 1/2?

zepdrix (zepdrix):

For the two blanks? Yes, good, in that order.

OpenStudy (anonymous):

okay thank you but how come the 1/2 didnt change?

zepdrix (zepdrix):

So you're probably thinking, well the angle changed from 30 to 60, why didn't the 1/2 change? But notice that the function also changed from sine to cosine. So two things changed, and it ended up giving us the same result. It'd be difficult to explain co-functions without going into some detail and drawing a triangle and all that...

OpenStudy (anonymous):

okay well thank you so much